Exercise 9-7
Job Cost for a Client

The partners of Mason and Hutton, attorneys-at-law, decide to implement an activity-based costing system for their firm. They identify the following three cost pools and budgeted amounts for each for the coming year: fringe benefits, $450,000; paralegal support, $250,000; and research support, $650,000. It is determined that the best cost driver for fringe benefits is professional labor dollars ($1,460,000); paralegal support is partner labor hours (4,000); and research support is professional labor hours (20,000).

A client, Jim Adams, requires 10 partner labor hours and 25 professional associate hours from Mason and Hutton, the law firm. Partners are paid $125 per hour, and associates make $60 per hour.

Compute the cost of the Adams' job. Round your interim calculations to two decimal places and your final answer to the nearest cent.

Its not 4652.50, or 4649.8
Answer #1

Ans.) Calculation of budgeted overhead rates
Fringe Benefits= 450000/1460000 = 0.308
Paralegal Support = 250000/4000 = 62.5
Research Support = 650000/20000 = 32.5

Calculation of Cost of Adam's Job = (10*125)+(25*60)= 2750
Fringe Benefit = 2750*0.308 = 847
Paralegal Support = 10 hrs * 62.5 = 625
Research Support = 25 hrs * 32.5 = 812.5
Total Cost of JOB = 2750+847+625+812.5 = $5034.5
(Is this the correct answer?)